In Stern-Gerlach experiment, calculate the force being exerted on the silver atoms due to the inhomogeneous magnetic field and hence calculate the maximum displacement occurs between the traces. 



1
Expert's answer
2021-01-11T11:34:49-0500

Answer

The inhomogenous of the magnetic field means that the atoms experience a sideways force given by

"F_z=-\frac{d U}{dz}"

Where U = −µ · B = −µzB is the potential energy of the silver atom in the magneic field becomes as

"F_z=\mu_z\frac{d U}{dz}"

Now,

maximum displacement occurs between the trace is given by

"Z=F_z(\frac{(d_1(d_1/2+d_2)) }{MV^2})"

Diagram can be shown as